OpenStudy (anonymous):

Solve: −3(x − 2) = 15

OpenStudy (anonymous):

divide both sides by -3

OpenStudy (anonymous):

Order of Operations= Start by dividing by -3. Then add 2 to both sides.

OpenStudy (mary.rojas):



OpenStudy (mary.rojas):

that is how you solve for x

OpenStudy (goldphenoix):

Bravo, mary.rojas. You want to do order of operation. Parenthesis comes before division. So you do ( x - 2) Now you can't subtract them because you have a variable. So you would want to do the distributive property. You distribute -3 to x and - 2. -3 * x = -3x -3 * -2 = 6. -3(x-2) become -3x + 6. Now your equation would look like: -3x + 6 = 15. You want to isolate the variable. To isolate the variable x, you want to subtract 6 from both side. 6 - 6 = 0 15 - 6 = 9. Now your equation read: -3x = 9 The final step is to divide both side by -3. -3 / -3 = 1. 9 / -3 = -3. Therfore, x = -3.

OpenStudy (mary.rojas):

:)